The price for a good A has risen from 175 rub. to 210 rub. The demand for a good B has increased from 5400 units to 7100 units. Calculate the cross- price elasticity of demand?


1
Expert's answer
2021-09-26T20:32:15-0400

Cross-price elasticity of demand"=" "\frac {percentage change in quantity of B} {percentage change in price of A}"

Change in quantity of B:

(7100"-" 5400)/ 5400"=" 31.48%

Change in price:

"(" 210"-" 175")" / 175"=" 20%

Cross price elasticity of demand"=" "\frac {31.48}{20}"


"=" 1.574